OpenStudy (mrnebula):

A 54-kg jogger is running at a rate of 3 m/s. What is the kinetic energy of the jogger? *18 J *81 J *162 J *243 J

OpenStudy (frostbite):

You know the formula for kinetic energy?

OpenStudy (mrnebula):

that y im asking i forgot

OpenStudy (frostbite):

Alright it is defined as: \[\Large E_{k}=\frac{ 1 }{ 2 }mv ^{2}\]

OpenStudy (mrnebula):

do u know?

OpenStudy (mrnebula):

whats E

OpenStudy (frostbite):

the energy.

OpenStudy (frostbite):

m is the mass and v the velocity

OpenStudy (mrnebula):

so i divide 54 by 1/2?

OpenStudy (mrnebula):

TEACH ME PLZ

OpenStudy (frostbite):

No, use the formula: \[\large E_{k}=\frac{ 1 }{ 2 } \times 54 kg \times (3 \frac{ m }{ s })^{2}\] Try evaluate it.

OpenStudy (mrnebula):

ok

OpenStudy (frostbite):

What do you get as result?

OpenStudy (mrnebula):

7.5

OpenStudy (frostbite):

That don't sound right. What did you press on your calc?

OpenStudy (mrnebula):

i used 0.5 as 1/2 and times it by 54 and times it by 3

OpenStudy (frostbite):

did you square 3?

OpenStudy (mrnebula):

i did that just now and i got 46

OpenStudy (frostbite):

Hmmmm sounds like you press something wrong. Here is what I got: http://lmgtfy.com/?q=1%2F2+*+54+kg+*+(3+m%2Fs)%5E2

OpenStudy (mrnebula):

yea i dun goofs but thinx atleast i have the formula

OpenStudy (frostbite):

No problem.